# Mapping the Spiral of Silence: Surveying Unspoken Opinions in Online Communities

### v1

@@ Appendix B Robustness Checks
WTSC $-0.37$ $0.08$ $0.00^{***}$ $-0.38$ $0.08$ $0.00^{***}$
Agreement Intensity $0.35$ $0.08$ $0.00^{***}$ $0.38$ $0.08$ $0.00^{***}$
Male (0/1) $-0.50$ $0.18$ $0.01^{**}$ $-0.52$ $0.18$ $0.00^{**}$
White (0/1) $0.03$ $0.17$ $0.85$ $0.06$ $0.17$ $0.74$
Democrat (0/1) $-0.64$ $0.17$ $0.00^{**}$ $-0.62$ $0.17$ $0.00^{**}$
Account Tenure $0.20$ $0.08$ $0.01^{*}$ $0.20$ $0.08$ $0.01^{*}$
Posting $0.40$ $0.08$ $0.00^{***}$ $0.41$ $0.08$ $0.00^{***}$
@@
Incongruency (0/1) $-0.73$ $0.18$ $0.00^{**}$ $-0.61$ $0.19$ $0.00^{**}$
Inclusion $0.38$ $0.08$ $0.00^{**}$ $0.28$ $0.10$ $0.00^{**}$
Diversity $0.06$ $0.08$ $0.54$ $0.05$ $0.10$ $0.60$
ModRatio $0.37$ $0.18$ $0.05^{*}$ $0.37$ $0.18$ $0.04^{*}$
Incongruency $\times$ Inclusion $0.36$ $0.18$ $0.05^{*}$
Incongruency $\times$ Diversity $0.01$ $0.18$ $0.95$
Incongruency $\times$ ModRatio $-0.15$ $0.21$ $0.47$
@@
Marginal $R^{2}$ $0.22$ $0.23$
Conditional $R^{2}$ $0.35$ $0.35$
Note: p$<$0.05; ∗∗p$<$0.01; ∗∗∗p$<$0.001
Table 10: The results of a linear mixed-effect model predicting a participant’s likelihood of sharing their opinion in a subreddit. We present results including a quadratic term for Content Removal Rate as an independent variable.
@@
WTSC $-0.37$ $0.08$ $0.00^{***}$
Agreement Intensity $0.35$ $0.08$ $0.00^{***}$
Male (0/1) $-0.49$ $0.18$ $0.01^{**}$
White (0/1) $0.04$ $0.17$ $0.82$
Democrat (0/1) $-0.65$ $0.17$ $0.00^{**}$
Account Tenure $0.20$ $0.08$ $0.01^{*}$
Posting $0.40$ $0.08$ $0.00^{***}$
@@
Incongruency (0/1) $-0.74$ $0.18$ $0.00^{**}$
Inclusion $0.39$ $0.08$ $0.00^{**}$
Diversity $0.06$ $0.08$ $0.51$
Content Removal Rate $1.31$ $0.69$ $0.06$
$\text{Content Removal Rate}^{2}$ $-1.10$ $0.72$ $0.13$
@@
Marginal $R^{2}$ $0.23$
Conditional $R^{2}$ $0.33$
Note: p$<$0.05; ∗∗p$<$0.01; ∗∗∗p$<$0.001
@@ Appendix C Deviations from Pre-registration
We pre-register our analyses on OSF. [^2] Pre-registration materials at https://osf.io/gze65?view_only=03f45025105a42b1b50cd6936ece06d0. We also report the following deviations from our pre-registration. First, we split community values into two separate hypotheses as we found perceived inclusion and diversity were measuring distinct constructs. We also updated the wording on our hypotheses to mention incongruent viewpoints. Second, we changed how we measured Content Removal Rate. In the pre-registration, we had planned to use 500 posts sampled from each subreddit; however, during analysis, we found this method underreported the amount of content removal compared to results from prior work (Jhaver, Bruckman, and Gilbert 2019). Thus, we decided to rely on Pushshift data (Baumgartner et al. 2020), which gave us access to a larger corpus of posts. Third, we remove the variable Mod Ratio from our models since it was correlated with Content Removal Rate. For robustness, we replicate our results using Mod Ratio as our measure for community moderation, finding similar results. We also added a random effect for subreddit. Finally, we conducted additional analyses using Upvote Likelihood as our dependent variable.
@@ Appendix D Other Techniques for Avoiding Self-Silencing
In our survey we also ask participants what other actions they are likely to take for viewpoints they would not share under their main account. As shown in Table 11, participants are most likely to “lurk,” or read discussion on a topic but not share their viewpoint. After lurking, the second most common option is to discuss the viewpoint offline with friends and family. Prior work (Hampton et al. 2014) found that people are more likely to engage in conversation on controversial political topics in face-to-face settings rather than on social media platforms.
We also find that participants leverage the affordances on Reddit to find other means for expressing their viewpoint. For example, sharing the viewpoint using a “throwaway account” (Leavitt 2015), a temporary account created under a different pseudonym than the user’s main account, is selected as a preferred option for 11 viewpoints. This suggests that in some cases, having a greater degree of anonymity, which a throwaway account offers, may help counteract silencing effects. Another option is to post the viewpoint to a different subreddit. When participants were asked why they decided to post to alternative subreddits, reasons included selecting subreddits that were smaller in size, were more related to the topic, or were more likely to agree with the viewpoint.
Table 11: Summary of actions that participants would take if they did not feel comfortable sharing a viewpoint, meaning they agreed with a viewpoint but reported that they were unlikely to share the opinion to a given subreddit.
Action N
Read discussion on a topic but not share viewpoint 68
Discuss with friends and family offline 34
Choose to ignore discussion on topic 19
Post viewpoint using a throwaway account 11
Post to a different subreddit 3
Other 3

### v2

@@ B.3. Replacing measures for moderation
WTSC 0.30-0.30 $0.25$ $0.22$ 0.30-0.30 $0.25$ $0.22$
Agreement Intensity $0.40$ $0.06$ $0.00^{***}$ $0.41$ $0.06$ $0.00^{***}$
Male (0/1) 0.71-0.71 $0.55$ $0.20$ 0.83-0.83 $0.54$ $0.13$
White (0/1) $0.11$ $0.53$ $0.83$ $0.19$ $0.53$ $0.72$
Democrat (0/1) 0.75-0.75 $0.49$ $0.13$ 0.75-0.75 $0.49$ $0.13$
Account Tenure $0.10$ $0.24$ $0.67$ $0.12$ $0.24$ $0.63$
Posting $0.64$ $0.24$ $0.01^{**}$ $0.65$ $0.24$ $0.01^{**}$
@@
Incongruency (0/1) 0.74-0.74 $0.14$ $0.00^{***}$ 0.57-0.57 $0.15$ $0.00^{***}$
Inclusion $0.07$ $0.10$ $0.49$ $0.03$ $0.11$ $0.81$
Diversity $0.02$ $0.10$ $0.84$ 0.06-0.06 $0.10$ $0.56$
ModRatio 0.15-0.15 $0.10$ $0.14$ 0.07-0.07 $0.10$ $0.53$
Incongruency $\times$ Inclusion $0.21$ $0.14$ $0.14$
Incongruency $\times$ Diversity $0.38$ $0.15$ $0.01^{*}$
Incongruency $\times$ ModRatio 0.35-0.35 $0.16$ $0.02^{*}$
@@
Marginal $R^{2}$ $0.25$ $0.26$
Conditional $R^{2}$ $0.81$ $0.81$
Note: p<<0.05; ∗∗p<<0.01; ∗∗∗p<<0.001
We examine changes to the model when replacing Content Removal Rate with ModRatio as our measure of moderation. ModRatio is the ratio of the number of subscribers to number of moderators in a subreddit. In this case, a higher ModRatio would mean less active moderation as there are fewer moderators per subscribers. We apply a logarithmic transformation with base 2 after adding a start-value of 1 to ModRatio. As shown in Table 11, there is a significant negative interaction between Incongruency and ModRatio ($\beta=-0.35$, $p=0.02$), mirroring our results when using Content Removal Rates. The trends with our independent variables are similar to when we used Content Removal Rate.
B.4. Examining quadratic terms for moderation
Table 12. The results of a linear mixed-effect model predicting a participant’s likelihood of sharing their opinion in a subreddit. We present results including a quadratic term for Content Removal Rate as an independent variable.
@@ B.4. Examining quadratic terms for moderation
WTSC 0.30-0.30 $0.25$ $0.23$
Agreement Intensity $0.40$ $0.06$ $0.00^{***}$
Male (0/1) 0.69-0.69 $0.55$ $0.22$
White (0/1) $0.10$ $0.53$ $0.86$
Democrat (0/1) 0.74-0.74 $0.49$ $0.14$
Account Tenure $0.08$ $0.24$ $0.75$
Posting $0.71$ $0.24$ $0.01^{**}$
@@
Incongruency (0/1) 1.09-1.09 $0.22$ $0.00^{***}$
Content Removal Rate $0.06$ $0.12$ $0.62$
$\text{Content Removal Rate}^{2}$ $0.04$ $0.11$ $0.70$
Incongruency $\times\text{Content Removal Rate}$ 0.02-0.02 $0.19$ $0.90$
Incongruency $\times\text{Content Removal Rate}^{2}$ $0.32$ $0.18$ $0.07$
@@
Marginal $R^{2}$ $0.25$
Conditional $R^{2}$ $0.81$
Note: p<<0.05; ∗∗p<<0.01; ∗∗∗p<<0.001
We also report results of our linear mixed-effects model after including a quadratic term for Content Removal Rate in Table 12. Again, we see similar trends between Incongruency and Share Likelihood. There is no significant relationship between the quadratic term and our dependent variable.
B.5. Treating sharing likelihood as ordinal
In Table 13, we use an ordinal mixed-effect regression to predict Share Likelihood, which is a Likert-scale with values ranging from $1$ to $7$. In the main body of the paper, we report results treating Share Likelihood as continuous, as is consistent with other works measuring the spiral of silence (Gearhart and Zhang, 2014, 2015; Willnat et al., 2002; Ho et al., 2013). This decision is also consistent with prior work in statistics, demonstrating the appropriateness of such an approach (Sullivan and Artino Jr, 2013; Norman, 2010).
Our conclusions about direction are robust across continuous and ordinal specifications; however, statistical significance varies with modeling assumptions. First, we note there is a statistically significant negative relationship between Incongruency ($\beta=-0.87$, $p<0.001$ in Model 3b) and the likelihood of sharing a viewpoint. Furthermore, we note a positive interaction effect between Diversity and opinion sharing ($\beta=0.26$) as well as a negative interaction effect between Content Removal Rate and opinion sharing ($\beta=-0.28$) although the interaction effects are not significant at our level of statistical power ($p=0.11$ for both).
Table 13. We present results using an ordinal mixed-effect model. The direction of the relationships are robust to the model reported in Sec. 5.
Model 3a Model 3b Model 3c
Coef. SE $p$ Coef. SE $p$ Coef. SE $p$
Dependent Variable
Share Likelihood $\leq$ 1 2.45-2.45 $0.68$ $0.00^{***}$ 2.84-2.84 $0.71$ $0.00^{***}$ 2.96-2.96 $0.71$ $0.00^{***}$
Share Likelihood $\leq$ 2 1.20-1.20 $0.68$ $0.08$ 1.53-1.53 $0.70$ $0.03^{*}$ 1.61-1.61 $0.70$ $0.02^{*}$
Share Likelihood $\leq$ 3 0.72-0.72 $0.68$ $0.29$ 1.02-1.02 $0.70$ $0.14$ 1.10-1.10 $0.70$ $0.00^{***}$
Share Likelihood $\leq$ 4 0.17-0.17 $0.68$ $0.81$ 0.46-0.46 $0.69$ $0.51$ 0.47-0.47 0.52-0.52 $0.46$
Share Likelihood $\leq$ 5 $0.83$ $0.68$ $0.22$ $0.62$ $0.69$ $0.37$ $0.59$ $0.70$ $0.40$
Share Likelihood $\leq$ 6 $1.97$ $0.68$ $0.00^{**}$ $1.83$ $0.70$ $0.00^{***}$ $1.80$ $0.70$ $0.01^{**}$
Controls
WTSC 0.42-0.42 $0.29$ $0.15$ 0.46-0.46 $0.29$ $0.12$ 0.47-0.47 $0.29$ $0.11$
Agreement Intensity $0.60$ $0.08$ $0.00^{***}$ $0.53$ $0.08$ $0.00^{***}$ $0.57$ $0.08$ $0.00^{***}$
Male (0/1) 0.58-0.58 $0.63$ $0.35$ 0.66-0.66 $0.64$ $0.30$ 0.70-0.70 $0.64$ $0.27$
White (0/1) $0.10$ $0.61$ $0.87$ $0.11$ $0.62$ $0.86$ $0.11$ $0.62$ $0.86$
Democrat (0/1) 0.77-0.77 $0.57$ $0.18$ 0.87-0.87 $0.58$ $0.13$ 0.89-0.89 $0.58$ $0.13$
Account Tenure $0.06$ $0.27$ $0.81$ $0.05$ $0.28$ $0.84$ $0.07$ $0.28$ $0.81$
Posting $0.80$ $0.28$ $0.00^{**}$ $0.77$ $0.29$ $0.01^{**}$ $0.80$ $0.29$ $0.01^{**}$
Independent Variables
Incongruency (0/1) 0.87-0.87 $0.17$ $0.00^{***}$ 0.80-0.80 $0.17$ $0.00^{***}$
Inclusion $0.10$ $0.12$ $0.41$ 0.02-0.02 $0.13$ $0.88$
Diversity $0.02$ $0.11$ $0.88$ 0.04-0.04 $0.12$ $0.76$
Content Removal Rate 0.07-0.07 $0.10$ $0.50$ $0.00$ $0.10$ $0.99$
Incongruency $\times$ Inclusion $0.49$ $0.19$ $0.01^{*}$
Incongruency $\times$ Diversity $0.26$ $0.17$ $0.11$
Incongruency $\times$ Content Removal Rate 0.28-0.28 $0.17$ $0.11$
Note: p<<0.05; ∗∗p<<0.01; ∗∗∗p<<0.001
@@ Appendix C Deviations from Pre-registration
We pre-register our analyses on OSF. We also report the following deviations from our pre-registration. First, we change our statistical model to include a crossed random effect between subreddit and participant, as we are capturing multiple observations per subreddit per participant. We updated our power analysis with this updated model to reach our target sample size of 270 responses. Second, we split community values into two separate hypotheses, as we found perceived inclusion and diversity were measuring distinct constructs. We also updated the wording on our hypotheses to mention incongruent viewpoints. Third, we changed how we measured Content Removal Rate. In the pre-registration, we had planned to use 500 posts sampled from each subreddit; however, during analysis, we found this method underreported the amount of content removal compared to results from prior work (Jhaver et al., 2019b). Thus, we decided to rely on Pushshift data (Baumgartner et al., 2020), which gave us access to a larger corpus of posts. Fourth, we remove the variable Mod Ratio from our models since it was correlated with Content Removal Rate. For robustness, we replicate our results using Mod Ratio as our measure for community moderation, finding similar results. Finally, we conducted additional analyses using Upvote Likelihood as our dependent variable, which are presented as post-hoc analyses.
@@ Appendix D Other Techniques for Avoiding Self-Silencing
In our survey, we also ask participants what other actions they are likely to take for viewpoints they would not share under their main account. Participants (N=68) are most likely to “lurk,” or read discussion on a topic but not share their viewpoint. After lurking, the second most common option is to discuss the viewpoint offline with friends and family (N=34). Prior work (Hampton et al., 2014) found that people are more likely to engage in conversation on controversial political topics in face-to-face settings rather than on social media platforms.